OpenStudy (anonymous):

What is the most reasonable estimate of 10.9 – 3.2? 9 8 7 6 @Melodysim @melanielovee4

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Thank you for helping me!

OpenStudy (mathstudent55):

I think the answer is 8. 10.9 is approximately 11. 3.2 is approximately 3. 11 - 3 = 8, so the estimate is 8, not 7.

OpenStudy (anonymous):

I disagree. But we're allowed our opinions.

OpenStudy (mathstudent55):

@melanielovee4 I agree with you that an estimate is just that, an estimate and not an exact answer, so different people may estimate it differently. Notice, though, that if you do the exact subtraction, 10.9 - 3.2, the answer is 7.7, which rounds off to 8. If your method is to simply consider 10.9 to be approximately 10 and 3.2 to be approximately 3, and then you estimate by subtracting 10 - 3 = 7, then what would you do if the numbers were 10.1 - 8.9? If you just do 10 - 8, you get 2, but if you round the numbers, 10.1 is approximately 10. 8.9 is approximately 9, so the estimate for 10.1 - 8.9 should be 1, not 2. In fact, 10.1 - 8.9 = 1.2 which is much closer to 1 than to 2.